# On combinatorial invariance of parabolic Kazhdan–Lusztig polynomials

### v2

@@ Abstract.
CG is supported by a Klarman Postdoctoral Fellowship at Cornell University.
@@ 1. Introduction
We say a Bruhat interval $[u,v]$ is a short edge interval if all edges $y\to y^{\prime}$ in the Bruhat graph restricted to $[u,v]$ have $\ell(y^{\prime})-\ell(y)=1$. Conjecture 1.1 was proven for short edge intervals by Brenti [6].
@@
Section 2 gives background on Bruhat order and Kazhdan–Lusztig polynomials. In Section 3 we introduce invariant collections of Bruhat intervals and prove Theorem 3.3 which gives general criteria the for combinatorial invariance of parabolic Kazhdan–Lusztig polynomials. Theorems 1.6, 1.7, 1.8 and 1.9 will be shown in Section 4 to follow from Theorem 3.3 and known special cases of Conjecture 1.1.
@@ 3. Invariant collections of intervals
We first prove a key lemma.
@@
Suppose $y=uw\in uW_{J}\cap[e,v]$. Since $u\in W^{J}$, we have $u<uw$, so $y\in[u,v]$. If we had $y\geq a$ for some $a\in A^{J}_{u,v}$, then we would have $y^{J}\geq a^{J}=a>u$ by Proposition 2.1, but this contradicts the uniqueness of the parabolic decomposition $y=uw$. Thus $y$ belongs to the right hand side of (1).
@@
    We have $R_{y_{1},v}=R_{\varphi(y_{1}),v_{2}}$ by upper $R$-invariance and hence by (2) we conclude

### v3

@@ 1. Introduction
We say a Bruhat interval $[u,v]$ is a short edge interval if all edges $y\to y^{\prime}$ in the Bruhat graph restricted to $[u,v]$ have $\ell(y^{\prime})-\ell(y)=1$. Conjecture 1.1 was proven for short edge intervals by Brenti [6]. Note that short edge intervals coincide with the $S_{3}$-free intervals studied there, by [14, Proposition 3.3].
@@
Section 2 gives background on Bruhat order and Kazhdan–Lusztig polynomials. In Section 3 we introduce invariant collections of Bruhat intervals and prove Theorem 3.3 which gives general criteria for the combinatorial invariance of parabolic Kazhdan–Lusztig polynomials. Theorems 1.6, 1.7, 1.8 and 1.9 will be shown in Section 4 to follow from Theorem 3.3 and known special cases of Conjecture 1.1.
@@ 3. Invariant collections of intervals
We first prove a key lemma. We remark that this lemma is also a special case of [21, Proposition 3.4], using the fact that $w\mapsto w^{J}$ is a projection in the sense described there [20].
@@
Suppose $y=uw\in uW_{J}\cap[e,v]$. Since $u\in W^{J}$, we have $u\leq uw$, so $y\in[u,v]$. If we had $y\geq a$ for some $a\in A^{J}_{u,v}$, then we would have $y^{J}\geq a^{J}=a>u$ by Proposition 2.1, but this contradicts the uniqueness of the parabolic decomposition $y=uw$. Thus $y$ belongs to the right hand side of (1).
@@
    We have $R_{y_{1},v_{1}}=R_{\varphi(y_{1}),v_{2}}$ by upper $R$-invariance and hence by (2) we conclude